# Learning local equivariant representations for quantum operators

### v3

@@ Abstract
Predicting quantum operator matrices such as Hamiltonian, overlap, and density matrices in the density functional theory (DFT) framework is crucial for understanding material properties. Current methods often focus on individual operators and struggle with efficiency and scalability for large systems. Here we introduce a novel deep learning model, SLEM (strictly localized equivariant message-passing) for predicting multiple quantum operators, that achieves state-of-the-art accuracy while dramatically improving computational efficiency. SLEM’s key innovation is its strict locality-based design, constructing local, equivariant representations for quantum tensors while preserving physical symmetries. This enables complex many-body dependence without expanding the effective receptive field, leading to superior data efficiency and transferability. Using an innovative SO(2) convolution technique, SLEM reduces the computational complexity of high-order tensor products and is therefore capable of handling systems requiring the $f$ and $g$ orbitals in their basis sets. We demonstrate SLEM’s capabilities across diverse 2D and 3D materials, achieving high accuracy even with limited training data. SLEM’s design facilitates efficient parallelization, potentially extending DFT simulations to systems with device-level sizes, opening new possibilities for large-scale quantum simulations and high-throughput materials discovery.
1.   Introduction
@@
Quantum operators, representing observables and the evolution of quantum systems, are the cornerstone of describing the microscopic world. In modern quantum science, the advent of density functional theory (DFT) [1, 2] has elevated single-particle quantum operators, such as the Kohn-Sham Hamiltonian, density matrix, and overlap matrix, to paramount importance in solving complex problems [3]. These operators play a crucial role in unraveling electronic structures, predicting material properties, and advancing quantum technologies. However, as we tackle larger and more complex systems, the efficient and accurate representation of these fundamental operators has emerged as a pressing challenge in computation, demanding new avenues for innovative methodologies.
Recent advances have incorporated machine learning techniques to accelerate DFT calculations by directly predicting DFT’s output of quantum operators, including charge density [4], overlapping matrix [5, 4], self-energy [6], wave function [4], and Hamiltonian matrix [7, 5, 8, 9, 4, 10].
Figure 1: Illustration of the local design of SLEM in comparison with conventional MPNN on a 1D graph. (a) MPNN information aggregation. (b) SLEM information aggregation. Balls and sticks represent nodes and edges; arrows show aggregation direction. $r_{\text{cut}}$: pre-defined cut-off; $r_{\text{eff}}$: effective cut-off after two hidden layer updates. $L$ is the layer index.
@@
By circumventing self-consistent DFT calculations, such methods have the potential to scale up the electronic structure calculations. Some of the approaches utilize Gaussian regressions [9], kernel-based models [9], and neural networks [11] to predict invariant Hamiltonian matrix blocks on localized frames. Notably, powerful equivariant message-passing neural networks (E-MPNNs) have demonstrated remarkable accuracy [12, 13, 14, 15, 16, 17, 18, 19, 20]. These networks ensure the equivariance of output tensor blocks, respecting the physical priors of atomic systems. Typically, they initialize the node and edge features with atomic species and the weighted projection of the spherical harmonics coefficient, then use iterative two-body updates to build many-body interactions. Despite their high accuracy, these iterative updates invariably enlarge the receptive field of node and edge features, incorporating numerous effective neighbours. This expansion limits parallelization and, consequently, the model’s scalability. The storage-intensive nature of quantum tensor prediction tasks exacerbates these limitations, as high-dimensional features in edges and nodes necessitate substantial GPU memory during training and prediction. Consequently, training such models on large datasets or predicting quantum operators for extensive atomic structures poses considerable challenges.
@@
Fortunately, the electrostatic screening counteracts the long-term dependency in a lot of material systems [21, 22, 23]. For example, for bulk Silicon, Ge, a typical screening length is around 5 a.u. [23]. Therefore, the predicted quantum operators can be decomposed into elements with localized dependency on atomic structures. Naturally, a strictly local model that avoids expanding the receptive field, while maintaining sufficient accuracy is at most preferred. In the field of machine-learning interatomic potentials (MLIPs), this concept has been applied in the Allegro model [13], where the strictly local representation of interatomic potentials achieves high accuracy comparable to E-MPNNs and is highly parallelizable, particularly for large systems. However, unlike modelling interatomic potential, which only concerns scalars (energy) and vectors (forces) (corresponding to angular momentum $l=0$ and $1$) from node features, the prediction of the general representation of quantum operators necessitates targeting both node and edge features on high-order spherical tensors (even up to $l=6$ or $8$). This adds complexity to constructing strictly localized models for node and edge updates. Another significant challenge in high-order tensor prediction tasks is computational complexity. To equivariantly mix the features of different angular momentum $l$, tensor products that scale as $\mathcal{O}$($l^{6}$) are required [19]. Achieving high accuracy often necessitates up to hundreds of thousands of training steps, which significantly limits the applicability of current models to heavy elements. This limitation hinders the development of a unified machine-learning DFT model that generalizes across the periodic table. Therefore, a crucial question, whether one can employ the strictly localized scheme to develop powerful equivariant atom and bond features that can efficiently and accurately reproduce the corresponding quantum tensor block elements with reduced computational complexity, naturally arises.
@@
This work presents a novel method, the strictly localized equivariant message-passing model (SLEM), implemented in the framework of DeePTB package [24] for efficient representations of quantum operators. SLEM employs a fully localized scheme to construct high-order node and edge equivariant features for a general representation of quantum operators, including the Hamiltonian and density matrix. As illustrated in Fig.1, the model embeds localized edge hidden states and utilizes them to construct localized node and edge features without including distant neighbours beyond a prefixed cutoff range $r_{\text{cut}}$. This design enables the SLEM model to generalize more effectively, parallelize more easily, and scale to larger systems. Additionally, a fast and efficient SO(2) convolution [19] is implemented to reduce the $\mathcal{O}$($l^{6}$) complexity, with edge-specific training weights, thereby further enhancing the model’s accuracy. As for the overlap matrix, it is typically required along with other operators for property calculations. Previous works have used another E-MPNN that doubles the size of the networks [10] or extracts it from DFT calculations [8, 5], incurring additional costs or incorporating out-of-loop computation steps that complicate inference. In contrast, we utilized the two-center integrals nature and parameterized the overlap matrix with spherical independent scalars (i.e., Slater-Koster-like parameters [25]). This method, inspired by the early work of DeePTB [24], fits the overlap operator representation alongside other operators with minimal additional cost.
This paper is organized as follows. In section 2, we review the related works, focusing on message-passing neural networks and equivariant message passing. In section 3, we detail the architecture of the SLEM model. Section 4 presents our experimental results, comparing the accuracy and efficiency of our model with existing models in reported data sets, and conducting generalization experiments. Finally, we summarize our findings and discuss future directions.
2.   Related works
@@
Message-passing neural networks (MPNNs) [26] have been widely applied in the modelling of atomic systems due to their exceptional accuracy in capturing the intricate relationships between atomic environments and physical properties [27]. In the message-passing scheme, atoms are treated as nodes in graphs, with bonds to neighbouring atoms represented as connected edges within a specified cutoff radius. The embedded atomic features are processed by trainable functions, generating messages from each edge to update the embeddings of central atoms. Formally, the MPNN framework can be summarized as follows:
Here, $\mathbf{e}^{ij,L}$ represents the edge features, $\mathbf{m}^{ij,L}$ denotes the messages, and $\mathbf{n}^{i,L}$ indicates the node features at layer $L$. $\mathcal{N}_{L}$, $M_{L}$, and $U_{L}$ are the trainable functions for the edge, message, and node updates. $\mathcal{N}(i)=\{j|r_{ij}<r_{\text{cut}}\}$ indicates all the neighbour atoms for atom $i$, with $r_{\text{cut}}$ being the predefined cutoff. This updating framework allows for the construction of many-body interactions and long-term dependencies, leading to strong performance across various applications. Previous works have predominantly utilized this updating scheme, achieving remarkably high precision in reported systems [28, 29, 12]. However, as updates progress, the effective cutoff radius ($r_{\text{eff}}=N\times r_{\text{cut}}$) for each atom’s features grows linearly with the number of update steps ($N$), as shown in Fig.1. Consequently, the effective neighbour list scales cubically with $r_{\text{eff}}$, making parallelization intractable. This issue is particularly severe when fitting representations of quantum operators, as the $r_{\text{cut}}$ defined by the cutoff radius for the linear combination of atomic orbitals (LCAO) basis is often larger than that defined in other tasks. A localized approach to constructing many-body interactions without including distant atoms has been proposed. For example, the Allegro model [13] modifies the updating rules by incorporating a hidden atom-pair state that depends explicitly on the direction vector and the center atom embedding. While this framework works successfully for potential energy prediction, it requires further improvement for fitting more general edge and node features. Other local methods avoid iterative updates that enlarge the receptive field, instead creating manually designed descriptors of the local environment [30, 31, 32, 33, 34, 35]. These methods are generally local but are criticized for insufficiency in describing high-order interactions.
Figure 2: The design of the SLEM model. (a) the hierarchy structure of the model. Starts with the structure information such as atomic number $Z_{i}$, and the radial and spherical part of the shift vector $\bm{r}_{ij}$ and $\mathbf{Y}^{ij}_{c,l}$, the initialized hidden $\mathbf{x}^{ij}$, $\mathbf{V}^{ij}$, edge and node features $\mathbf{e}^{ij}$, $\mathbf{n}^{i}$ would be generated. The initialized two-body hidden features would be used to predict the Slater-Koster-like Overlap parameters directly and transformed to construct (off-)diagonal blocks for the overlap operator. Others then, will update iteratively using the design’s strict localized updating scheme. (b)-(d) shows the details of the hidden update (b), edge update (c), and node update (d). After the updates, the node and edge features would be used to construct the (off-)diagonal blocks using the Wigner-Eckart theorem for quantum operators such as density matrix and Hamiltonian.
@@
There are certain natures of the physical quantities that should be invariant or equivariant under e.g. the spatial and temporal operations. For example, under translation, rotation, and reflection, the scalars such as the energy of a system are invariant, while the tensors, such as forces and Hamiltonian matrices should transform accordingly. Such three kinds of operations constitute the group operations of E(3), known as the Euclidian Group (O(3) if without translation). To model such equivariant quantities, a set of neural network models has been developed that incorporate operations that preserve this equivariance. [36, 37, 38, 39] These neural networks possess physical priors, as their operations ensure the output transforms in sync with the input, making them more generalizable, accurate, and data-efficient in predicting physical quantities. Formally, an equivariant operation from vector space X to vector space Y is defined such that:
$f(D_{X}[g]x)=D_{Y}[g]f(x)\quad\forall g\in G,\forall x\in X$
where $D_{X}[g]\in GL(X)$ is the representation of group element $g$ on vector space $X$. Since the representation takes the form of the direct sum of irreducible representations of the group $G$, it suggests that the vectors $x$ and $y$ themselves are composed by the irreducible representation (irreps for short) of group $G$. Usually, we consider here for operations of group O(3)(since translational symmetry is naturally satisfied by using relative coordinates), the irreps are the spherical tensors with angular momentum index $l$ and parity $p$. So features $x,y$ should also be indexed by $l,p$. For each irrep with angular momentum $l$, an internal degree of freedom, denoted by $m$ exists such that $|m|\leq l$. Therefore, each irreps of order $l$ belongs to a subspace of dimension $12l+$ 1. Once the $l,p$ index is defined, one can construct the group operations, which are represented by the $(2l+1)\times(2l+1)$ Wigner-D matrices.
Performing numerical operations on such irreps features follows certain rules. For the tensors with the same $l$, addition and subtraction can be conducted. To mix up the tensors of different $l$ indices, an important operation called tensor product ($\otimes$) is applied, which is considered the generalized operation of multiplication on spherical tensors. More formally:
@@
Where $C_{(l_{1},m_{2})(l_{2},m_{2})}^{(l_{3},m_{3})}$ are Clebsch-Gordan (CG) coefficients. This operation is found extremely useful in constructing edge and node updating functions, as seen in []. However, the conventional tensor product has the time and memory scales of $O(l_{max}^{6})$ [19] where $l_{max}$ is the maximum angular momentum of $\mathbf{x}$ and $\mathbf{y}$. Such complexity limits the application like the equivariant forces fields where high inference speed is vital. It poses an even greater challenge for quantum tensor prediction tasks, which require tensor products on very high-order spherical tensors. For example, constructing blocks of $d$-$d$, $f$-$f$ and $g$-$g$ orbital pairs require irreps of maximum order $l=4$, $l=6$ and $l=8$, respectively. Such high costs make training for large-size systems nearly impossible.
3.   Model Architecure
III.1 Parameterize Equivariant Quantum Operators
The quantum operators $\hat{O}$, such as the Hamiltonian and density matrix in the LCAO-based DFT framework, follow the equivariance of the $E(3)$ group and can thus be expressed in tensorized features using group theory. This process is illustrated in Fig. 2. Specifically, the $E(3)$ group consists of spatial rotation, translation, and reflection. Translational symmetry is automatically satisfied by performing operations only on relative coordinates, while rotation and reflection require special treatments. Under the LCAO basis, the quantum tensor block can be expressed as:
$O_{l_{1},l_{2},m_{1},m_{2}}^{i,j}=\langle i,l_{1},m_{1}|\hat{O}|j,l_{2},m_{2}\rangle$
Here $i$ and $j$ denote atomic sites, while $l_{1},m_{1}$, and $l_{2},m_{2}$ refer to the atomic orbitals of site $i$ and site $j$ respectively. The terms $m_{1}$ and $m_{2}$ are the magnetic quantum numbers associated with the angular momentum $l_{1}$ and $l_{2}$. $O_{l_{1},l_{2},m_{1},m_{2}}^{i,j}$ are the matrix element of operator $\hat{O}$. When the coordinates of the system are transformed by a rotation operation $\mathbf{R}$, equivariance requires the quantum tensor to transform accordingly as:
$O^{\prime i,j}_{l_{1},l_{2},m_{1}^{\prime},m_{2}^{\prime}}=\sum_{m_{1},m_{2}}D(\mathbf{R})^{l_{1}}_{m_{1},m_{1}^{\prime}}D^{*}(\mathbf{R})^{l_{2}}_{m_{2},m_{2}^{\prime}}O_{l_{1},l_{2},m_{1},m_{2}}^{i,j}$ (1)
Where $O^{\prime}$ is the equivariant block after transformation, and $D(\mathbf{R})^{l}_{m,m^{\prime}}$ is the Wigner-D matrix of order $l$. To construct such equivariance blocks from irreps, the Wigner-Eckart theorem can be applied, which decomposes the operator indexed by $l_{1},l_{2}$ into a single index $l_{3}$ that satisfies $|l_{1}-l_{2}|\leq l_{3}\leq(l_{1}+l_{2})$, which states:
@@ III.1 Parameterize Equivariant Quantum Operators
Here, the edge ($i\neq j$) and node ($i=j$) features $o_{l_{3},m_{3}}^{i,j}$ are grouped by the index $m$ into vectors of $\mathbf{o}_{c,l}^{i,j}$ with $c$ accounting for multiple tensors for the same $l$. These features can be computed for hopping ($i\neq j$) and onsite ($i=j$) elements of the quantum operators. Further by leveraging the Hermitian nature of quantum operators, the parameterized elements can be further reduced to upper diagonal blocks.
The decomposed features $\mathbf{o}_{c,l}^{i,j}$ are aligned with the network prediction through a rescaling procedure. The target node/edge equivariant features $\mathbf{o}^{i,i/j}_{c,l}$ can be standarized and decomposed into constant norms and biases, and the normalized features that have balanced variance. Formally:
$\mathbf{o}^{i,i/j}_{c,l}$ $=\sigma^{Z_{i},Z_{i/j}}_{c,l}\hat{\mathbf{o}}^{i,i/j}_{c,l}+\mu^{Z_{i},Z_{i/j}}_{c,l}\delta_{l,0}$
Here, $Z_{i}$, $Z_{j}$ are the atom species of atom $i$ and $j$, $\sigma^{Z_{i},Z_{i/j}}_{c,l}$ and $\mu^{Z_{i},Z_{i/j}}_{c,l}$ are norm and bias value for each atom and atom-pair. These values are derived from the dataset statistics and subsequently applied as weights and shift biases in an atom/bond type-specific scaling layer. The improvement of the standardization procedure on enhancing model convergence is evident. For the tensor operators in LCAO basis, the diagonal blocks (or node features) often exhibit highly unbalanced norms across irreps of different $l$, hindering the model from learning a smooth function. Additionally, different irreps in the off-diagonal blocks (or edge features) demonstrate diverse decaying behaviors, making it challenging to learn these decays implicitly. By normalizing in this manner, the decay function for each irrep is constrained to a value ranging from 1 to 0. This facilitates the use of hidden states and a few layers of a Multilayer Perceptron (MLP) activated by a ReLU-like function, to explicitly project the decaying value from the preserved hidden scalar features for each edge. Detailed analysis of the decaying behavior of edge irreps and the norm of node irreps can be found in the supplementary materials (SM).
At this stage, the objective is to generate the node and edge features $\mathbf{n}^{i}_{c,l}$, $\mathbf{e}^{ij}_{c,l}$, and to supervise them towards achieving decomposed target features $\hat{\mathbf{o}}^{i,i}_{c,l}$ and $\hat{\mathbf{o}}^{i,j}_{c,l}$. Once a satisfactory accuracy is attained, the inversion of the Wigner-Eckart theorem facilitates the reconstruction of predicted operator representations such as Hamiltonian and density matrix blocks:
$O^{i,j}_{l_{1},l_{2},m_{1},m_{2}}=\sum_{l_{3},m_{3}}C_{(l_{1},m_{1})(l_{2},m_{2})}^{(l_{3},m_{3})}o_{l_{3},m_{3}}^{i,j}$
@@
III.2 Parameterize of Invariant Overlap Operators

### v4

@@ Abstract
Predicting quantum operator matrices such as Hamiltonian, overlap, and density matrices in the density functional theory (DFT) framework is crucial for material science. Current methods often focus on individual operators and struggle with efficiency and scalability for large systems. Here we introduce a novel deep learning model, SLEM (strictly localized equivariant message-passing) for predicting multiple quantum operators, that achieves state-of-the-art accuracy while dramatically improving computational efficiency. SLEM’s key innovation is its strict locality-based design for equivariant representations of quantum tensors while preserving physical symmetries. This enables complex many-body dependency without expanding the effective receptive field, leading to superior data efficiency and transferability. Using an innovative SO(2) convolution and invariant overlap parameterization, SLEM reduces the computational complexity of high-order tensor products and is therefore capable of handling systems requiring the $f$ and $g$ orbitals in their basis sets. We demonstrate SLEM’s capabilities across diverse 2D and 3D materials, achieving high accuracy even with limited training data. SLEM’s design facilitates efficient parallelization, potentially extending DFT simulations to systems with device-level sizes, opening new possibilities for large-scale quantum simulations and high-throughput materials discovery.
1 Introduction
@@
Quantum operators, representing observables and the evolution of quantum systems, are the cornerstone of describing the microscopic world. In modern quantum science, the advent of density functional theory (DFT) Hohenberg & Kohn (1964); Kohn & Sham (1965) has elevated single-particle quantum operators, such as the Kohn-Sham Hamiltonian, density matrix, and overlap matrix, to paramount importance in solving complex problems Jones (2015). These operators play a crucial role in unravelling electronic structures, predicting material properties, and advancing quantum technologies. However, as we tackle larger and more complex systems, these fundamental operators’ efficient and accurate representation has emerged as a pressing challenge in computation, demanding new avenues for innovative methodologies.
@@
Recent advances have incorporated machine learning (ML) techniques to accelerate DFT calculations by directly predicting DFT’s output of quantum operators, including charge density Unke et al. (2021), overlapping matrix Yu et al. (2023); Unke et al. (2021), self-energy Dong et al. (2024), wave function Unke et al. (2021), and Hamiltonian matrix Yin et al. (2024); Yu et al. (2023); Gong et al. (2023); Nigam et al. (2022); Unke et al. (2021); Zhong et al. (2023). By circumventing self-consistent DFT calculations, such methods have the potential to scale up the electronic structure calculations. Some of the approaches utilize Gaussian regressions Nigam et al. (2022), kernel-based models Nigam et al. (2022), and neural networks Li et al. (2022) to predict invariant Hamiltonian matrix blocks on localized frames. Notably, powerful equivariant message-passing neural networks (E-MPNNs) have demonstrated remarkable accuracy Han et al. (2024); Musaelian et al. (2023); Batatia et al. (2022); Joshi et al. (2023); Liao & Smidt (2022); Liao et al. (2023); Simeon & De Fabritiis (2024); Passaro & Zitnick (2023); Zitnick et al. (2022). These networks ensure the output tensor blocks’ equivariance, respecting atomic systems’ physical priors. Typically, they use iterative updates to build many-body interactions, achieving high accuracy while enlarging the receptive field. This limits parallelization and, consequently, the model’s scalability. The storage-intensive quantum tensor prediction tasks exacerbate these limitations, posing considerable challenges for training such models on large datasets or predicting quantum operators for extensive atomic structures .
Figure 1: Local design of SLEM vs MPNN on 1D graph.(a) MPNN aggregation. (b) SLEM aggregation. Balls: nodes, sticks: edges, arrows: aggregation direction. $r_{\text{cut}}$: predefined cutoff; $r_{\text{eff}}$: effective cutoff after 2 layer updates. $L$: layer index.
@@
Fortunately, the electrostatic screening counteracts the long-term dependency in a lot of material systems Huckel & Debye (1923); Resta (1977); Ninno et al. (2006). Therefore, quantum operators can be decomposed into elements dependent locally on atomic structures, which natrually prefers a strictly local model that avoids expanding the receptive field. The Allegro model Musaelian et al. (2023) applied this concept to build ML interatomic potentials (MLIPs), achieving high accuracy and parallelizability. While MLIPs only concern scalars (energy) and vectors (forces) (angular momentum $l=0$ and $1$) on each node, predicting quantum operators necessitates targeting both node and edge features on high-order spherical tensors (even up to $l=6$ or $8$). This requires locality and representability for both node and edge. Another significant challenge is computational complexity. To equivariantly mix the features of different angular momentum $l$, tensor products that scale as $\mathcal{O}$($l^{6}$) are required Passaro & Zitnick (2023). This makes model training, especially on heavy atoms, extremely slow. This limitation hinders the development of a unified ML DFT model that generalizes across the periodic table.
@@
This work presents a novel method, the strictly localized equivariant message-passing model (SLEM), for efficient representations of quantum operators. SLEM employs a fully localized scheme to construct high-order node and edge equivariant features for a general representation of quantum operators, including the Hamiltonian and density matrix. As illustrated in Fig. 1, the model embeds localized edge hidden states and utilize them to construct localized node and edge features without including distant neighbours beyond a fixed cutoff range $r_{\text{cut}}$. This design enables the SLEM model to generalize better, parallelize easier, and scale to larger systems. Additionally, a fast and efficient SO(2) convolution Passaro & Zitnick (2023) is implemented to reduce the $\mathcal{O}$($l^{6}$) complexity, with edge-specific training weights, thereby further enhancing the model’s accuracy. As for the overlap matrix, it is typically required for property calculations. Previous works have used another E-MPNN that doubles the network sizes Zhong et al. (2023) or extracts it from DFT calculations Gong et al. (2023); Yu et al. (2023), incurring additional costs or incorporating out-of-loop computation steps that complicate inference. In contrast, we utilized the two-centre integrals and parameterized the overlap matrix with spherical independent scalars (i.e., Slater-Koster (SK) parameters Slater & Koster (1954)). This method, inspired by the work of DeePTB Gu et al. (2023), fits the overlap operator representation with minimal additional cost.
2 Related works
@@
Message-passing neural networks (MPNNs) Gilmer et al. (2017) have been widely applied in the modelling of atomic systems due to their exceptional accuracy in capturing the intricate relationships between atomic environments and physical properties Schütt et al. (2020). Previous works have predominantly utilized this scheme, achieving remarkably high precision in reported systems Schütt et al. (2018); Satorras et al. (2021); Han et al. (2024). However, as the MPNNs updates, the effective cutoff radius ($r_{\text{eff}}=N\times r_{\text{cut}}$) for each atom’s features grows linearly with the number of update steps ($N$), as shown in Fig. 1. Consequently, the effective neighbour list scales cubically with $r_{\text{eff}}$, making parallelization intractable. Allegro Musaelian et al. (2023) achieved locality modifies the updating rules by incorporating a hidden atom-pair state that depends partially on the centre atom. While this framework works successfully for scalars and $l$=1 vectors in potential energy prediction, it requires further improvement for fitting more general edge and node features. Other local methods avoid iterative updates that enlarge the receptive field, instead creating manually designed descriptors of the local environment Behler & Parrinello (2007); Bartók et al. (2010); Thompson et al. (2015); Zhang et al. (2018a); Wang et al. (2018); Zhang et al. (2018b). These methods are generally local, but often balance locality and representation capacity.
@@
Physical quantities, under the law of nature, should be invariant or equivariant under the spatial and temporal symmetry operations. To model such quantities, a set of neural network models has been developed utilizing equivariant operations. Thomas et al. (2018); Weiler et al. (2018); Kondor et al. (2018); Kondor (2018) These neural networks possess physical priors to ensure outputs transform in sync with inputs, making them more generalizable, accurate, and data-efficient in predicting physical quantities. Formally, an equivariant operation from vector space $X$ to $Y$ is defined such that:
$f(D_{X}[g]\mathbf{x})=D_{Y}[g]f(\mathbf{x})\quad\forall g\in G,\forall\mathbf{x}\in X$
where $D_{X}[g]\in GL(X)$ is the representation of group element $g$ on vector space $X$. Here we consider $O(3)$ group, then $\mathbf{x},\mathbf{y}$ can be composed by irreducible representation (irreps for short) that are the spherical tensors with angular and magnetic momentum index $l,m$, and parity $p$ such that $|m|\leq l$. Irreps with the same $l$ support addition/subtraction, while a generalized multiplication is defined as tensor product ($\otimes$):
@@
Where $C_{(l_{1},m_{2})(l_{2},m_{2})}^{(l_{3},m_{3})}$ are Clebsch-Gordan (CG) coefficients. Conventional tensor product has the time and memory scales of $O(l_{\text{max}}^{6})$ Passaro & Zitnick (2023) where $l_{\text{max}}$ is the maximum angular momentum in $\mathbf{x}$ and $\mathbf{y}$. Such complexity poses great challenges for quantum tensor prediction. For example, constructing blocks of $f$-$f$ and $g$-$g$ orbital pairs require irreps of maximum order $l=6$ and $l=8$. Such high costs make training for large-size systems nearly impossible.
3 Model Architecure
Figure 2: Design of the SLEM model. (a) Hierarchical structure of the model. Starts with the atomic number $Z_{i}$, the radial and spherical part of the shift vector $\bm{r}_{ij}$ and $\mathbf{Y}^{ij}_{c,l}$, the initialized hidden features $\mathbf{x}^{ij}$, $\mathbf{V}^{ij}$, along with edge and node features $\mathbf{e}^{ij}$, $\mathbf{n}^{i}$, are generated. The two-body hidden features predict the SK parameters constructing (off-)diagonal blocks for the overlap operator. Others features are then iteratively updated using the designed strictly localized updating scheme. (b)-(d) shows the hidden update (b), edge update (c), and node update (d). Node and edge features are used to construct the diagonal blocks for quantum operators.
3.1 Parameterize Equivariant Quantum Operators
The equivariant parameterization of quantum operators $\hat{O}$, such as the Hamiltonian and density matrix in the LCAO-based DFT framework, is illustrated in Fig. 2. The matrix element of operator $\hat{O}$ can be expressed as:
$O_{l_{1},l_{2},m_{1},m_{2}}^{i,j}=\langle i,l_{1},m_{1}|\hat{O}|j,l_{2},m_{2}\rangle$ (1)
Here $i$ and $j$ denote atomic sites, while the angular and magnetic momentum index $l,m$ label the atomic orbitals of the site. We apply the Wigner-Eckart theorem to decomposes the operator indexed by $l_{1},l_{2}$ into a single index $l_{3}$ that satisfies $|l_{1}-l_{2}|\leq l_{3}\leq(l_{1}+l_{2})$:
@@ 3.1 Parameterize Equivariant Quantum Operators
Here, the edge ($i\neq j$) and node ($i=j$) features $o_{l_{3},m_{3}}^{i,j}$ are grouped by the index $m$ into vectors of $\mathbf{o}_{c,l}^{i,j}$ with $c$ accounting for multiple tensors for the same $l$. These features can be computed for hopping ($i\neq j$) and onsite ($i=j$) elements of the quantum operators. Further, by leveraging the Hermitian nature of quantum operators, the parameterized elements can be reduced to upper diagonal blocks. This is almost half the demanding storage. Then, we standardised $\mathbf{o}^{i,j}_{c,l}$ to balance the variance. Formally:
$\mathbf{o}^{i,j}_{c,l}=\sigma^{Z_{i},Z_{j}}_{c,l}\hat{\mathbf{o}}^{i,j}_{c,l}+\mu^{Z_{i},Z_{j}}_{c,l}\delta_{l,0}$ (3)
Here, $Z_{i}$, $Z_{j}$ are the atom species of atom $i$ and $j$, $\sigma^{Z_{i},Z_{j}}_{c,l}$ and $\mu^{Z_{i},Z_{j}}_{c,l}$ are norm and bias value for each atom and atom-pair. These values are derived from the dataset statistics and applied as weights and biases in an atom/bond type-specific scaling layer. This step helps to resolve the unbalanced norms across diagonal and off-diagonal elements of quantum operators while facilitating using a ReLU-activated network to learn the normalized radial dependent decaying function from 1 to 0.
After supervising on the normalized features $\hat{\mathbf{o}}^{i,j}_{c,l}$, inverse transform from Eq. 2 is applied to reconstruct the predicted operator representations such as Hamiltonian and density matrix blocks:
$O^{i,j}_{l_{1},l_{2},m_{1},m_{2}}=\sum_{l_{3},m_{3}}C_{(l_{1},m_{1})(l_{2},m_{2})}^{(l_{3},m_{3})}o_{l_{3},m_{3}}^{i,j}$ (4)
@@
3.2 Parameterize of Invariant Overlap Operators